OpenStudy (anonymous):

with a interest of 1,832.00 and a principal of 16,000 for 206 days, what is the rate using the ordinary interest method?

OpenStudy (anonymous):

A = total cost = interest + principal A = P(1+i)^n A/P = (1+i)^n n rt(A/P) = 1+i n rt(A/P) - 1= i

OpenStudy (anonymous):

if we have a daily interest rate of i, then we have a yearly rate of 360i 365i depending on what part of the world you are in

OpenStudy (jdoe0001):

what formula are you supposed to use again?